Which has maximum number of atoms?
a. 24 g of C (12)
b. 56 g of Fe (56)
c. 27 g of A1 (27)
d. 108 g of Ag (108)

Grade:11

6 Answers

Jitender Pal
9 years ago
Sol. Number of atoms = number of moles × Avogadro’s number (N A) ⇒ Number of atoms in 24 g C = 24/12×N A= 2N A Number of atoms in 56 g of Fe = 56/56N A = N A Number of atoms in 27 g of A1 = 27/27N A = N A Number of atoms in 108 g of Ag = 108/108N A = N A Hence, 24 g of carbon has the maximum number of atoms.
Anjali kumari
6 years ago
(A) As number of moles= given mass /molecular mass. or 24/12=2 moles.(B) number of moles=56/56=1 mole(C) number of moles=27/27=1 mole(D) number of moles=108/108=1 moles(A) will have the maximum number of atoms because it has 2 moles .As number of moles increases then the number of atoms will also increases.
pradnya
5 years ago
no of moles=given mass/molecular mass
C= 24/12= 2
Fe=56/56=1
Al =27/27=1
Ag=108/108=1
therefore , 24g of C has maximum number of atoms.
Anjali kumari
5 years ago
As in ( 1) the numbers of moles of carbon=2moles.And in (2) ,(3),(4) the number of moles =1moles.So maximum number of atoms=2 moles. Therefore the answer is (A).
Yashika
5 years ago
No.of moles= mass/molecular weight.In case 1 Molecular weight of C is 12 So 12/12=1 Molecular weight of Fe is 56 so 56/56=1 In the same way 27/27=1 and 108/108=108 So the correct option is (A) as it has 2 moles.
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Number of atoms = number of moles × Avogadro’s number (N A)
⇒ Number of atoms in 24 g C = 24/12×N A= 2N A
Number of atoms in 56 g of Fe = 56/56N A = N A
Number of atoms in 27 g of A1 = 27/27N A = N A
Number of atoms in 108 g of Ag = 108/108N A = N A
Hence, 24 g of carbon has the maximum number of atoms.

Thanks and Regards